OpenStudy (anonymous):

Manuel’s paintball team won $1200 in a local tournament. How much money, m, will each player get if the winnings, w, are divided evenly among 8 players, p?

OpenStudy (anonymous):

1200/p=m

OpenStudy (anonymous):

m=w/p m=1200/p m=1200/8 m=150

OpenStudy (anonymous):

1200/8=150